# A comparative high-resolution spectroscopic analysis of in situ and accreted globular clusters

### v1

@@
Globular clusters (GCs) are extremely intriguing systems that help in reconstructing the assembly of the Milky Way via the characterization of their chemo-chrono-dynamical properties. In this study, we use high-resolution spectroscopic archival data from UVES at VLT and UVES-FLAMES at VLT to compare the chemistry of GCs dynamically tagged as either Galactic (NGC 6218, NGC 6522 and NGC 6626) or accreted from distinct merging events (NGC 362 and NGC 1261 from Gaia-Sausage-Enceladus, Ruprecht 106 from the Helmi Streams) in the metallicity regime where abundance patterns of field stars with different origin effectively separate . We find remarkable similarities in the abundances of the two Gaia-Sausage-Enceladus GCs across all chemical elements. They both display depletion in the -elements (Mg, Si and Ca) and statistically significant differences in Zn and Eu compared to in situ GCs. Additionally, we confirm that Ruprecht 106 exhibits a completely different chemical makeup compared to the other target clusters, being underabundant in all chemical elements. This demonstrates that when high precision is achieved, the abundances of certain chemical elements not only can efficiently separate in situ from accreted GCs, but can also distinguish among GCs born in different progenitor galaxies. In the end, we investigate the possible origin of the chemical peculiarity of Ruprecht 106. Given that its abundances do not match the chemical patterns of field stars associated to its most likely parent galaxy (i.e., the Helmi Streams), especially being depleted in the -elements, we interpret Ruprecht 106 as originating in a less massive galaxy compared to the progenitor of the Helmi Streams.
@@
The latest data releases of the ESA/Gaia mission (Gaia Collaboration et al. 2021, 2023) have instigated a profound transformation in our comprehension of the early chronicles of the Milky Way (MW). This enormous 6-D phase space dataset led to a clearer understanding of the roles played by various mergers in shaping the Galactic halo (see Helmi 2020 for a review), as predicted by simulations in the CDM cosmological framework (White & Frenk 1991, Moore et al. 1999, Helmi & de Zeeuw 2000, Newton et al. 2018). The imprint of accretion events manifest dynamically as stellar streams (Helmi et al. 1999, Ibata et al. 2024), substructures discernible in phase space (e.g., Helmi et al. 2018, Belokurov et al. 2018, Myeong et al. 2018, Koppelman et al. 2019) and dwarf galaxies whose disruption is currently ongoing (Ibata et al. 1994, Majewski et al. 2003). During such events, not only field stars but also globular clusters (GCs) may survive the merging processes, contributing to the formation of the present-day MW GC system (Brodie & Strader 2006, Peñarrubia et al. 2009, Bellazzini et al. 2020, Trujillo-Gomez et al. 2021). Leveraging the extremely precise kinematic data provided by the Gaia mission, the orbits of MW GCs have been reconstructed with exceptional precision. Through the analysis of their orbital properties, MW GCs have been classified into distinct groups by several authors (Massari et al. 2019, Forbes 2020, Callingham et al. 2022, Chen & Gnedin 2024), distinguishing those accreted from disrupted dwarf galaxies, such as Gaia-Sausage-Enceladus (GSE, Belokurov et al. 2018, Helmi et al. 2018), Sequoia (Myeong et al. 2018) or the Helmi Streams (Helmi et al. 1999) from those formed in situ. Complications may arise in the interpretation of these findings due to the complex and overlapping distribution of GCs in orbital parameters spaces (Callingham et al. 2022). This complexity is further exacerbated when accounting for the impact of a non-static potential in the computation of a GC orbit (Amarante et al. 2022, Belokurov et al. 2023, Pagnini et al. 2023, Chen & Gnedin 2024). Nonetheless, it has been established that dynamically tagged groups of GCs lay on different sequences in the age-metallicity space (Kruijssen et al. 2019, Massari et al. 2019, Myeong et al. 2019), confirming the dual nature of the MW GC system in this plane already found in the pre-Gaia era (Marín-Franch et al. 2009, Forbes & Bridges 2010, Leaman et al. 2013, VandenBerg et al. 2013). In particular, it has been demonstrated that older systems at fixed metallicity are more likely to be born in galaxies with a higher star formation efficiency, like the MW in its earliest phases, compared to dwarf galaxies (e.g., Massari et al. 2023).
@@
This intricate perspective of the assembly history of the MW can be further enriched by the addition of chemical abundance data, obtained either from high-resolution spectra and from large spectroscopic surveys (Venn et al. 2004, Nissen & Schuster 2010, Helmi et al. 2018, Horta et al. 2020, Minelli et al. 2021, Limberg et al. 2022, Malhan et al. 2022, Naidu et al. 2022a, Horta et al. 2023, Monty et al. 2023a, Ceccarelli et al. 2024). However, when looking at the distribution of GCs in chemical spaces, the differences between these two populations are subtle, making the interpretation of the results not trivial. For instance, Recio-Blanco (2018) highlights a very low scatter in the ratio among GCs with regardless of their origin. On a similar note, Horta et al. (2020) employed the [Si/Fe] ratio from APOGEE DR17 (Abdurro’uf et al. 2022) of 46 GCs to explore whether their distribution in chemical spaces reflect the kinematically defined classification by Massari et al. (2019). Interestingly, they observed that the positions of both in-situ and accreted subgroups in this chemical space align with those of their field stars counterparts. However, they were unable to effectively identify different behaviours from GCs that were brought in by separate progenitors, possibly because of the higher precision needed. Recently, Belokurov & Kravtsov (2024) discriminated between Galactic and extragalactic origin for MW GCs using the [Al/Fe] ratios, building up on the findings by Belokurov & Kravtsov (2022). However, caution should be exercised when using the abundances of light elements in GCs, owing to the impact of multiple populations (Bastian & Lardo 2018). Finally, Monty et al. (2023a, b) showcased the efficacy of differential chemical analysis on GCs. Their works not only revealed chemical inhomogeneity at a precision level of 0.02 dex for all chemical elements in GCs, but also showed distinctions in the chemical compositions of two GCs, NGC 288 and NGC 362, associated with the same progenitor, GSE. Specifically, NGC 288 and NGC 362 exhibit dissimilarities in their chemical make-ups, particularly in neutron-capture elements. This difference could be interpreted either as stemming from chemical inhomogeneities in the progenitor or as a consequence of internal chemical evolution within the GCs. In light of all the mentioned complexities, which underline the challenges associated with interpreting results when comparing diverse studies, it becomes evident that maintaining homogeneity is imperative in conducting chemical analyses aimed at verifying the origin of GCs.
@@
In this study, we will derive fully homogeneous detailed elemental abundances for several species of six GCs with similar metallicity consistently associated with different progenitors (see below) by all post-Gaia studies (Massari et al. 2019, Forbes 2020, Callingham et al. 2022, Chen & Gnedin 2024). These GCs lie in the metallicity range best suitable for chemical tagging studies, where abundance trends of in situ and accreted objects start to significantly part ways (Nissen & Schuster 2010, Helmi et al. 2018, Matsuno et al. 2022b). Three of the target clusters, namely NGC 362, NGC 1261, and Ruprecht 106 (Rup106 hereafter), lay on the younger branch of the Galactic GCs age-metallicity relation (AMR, Dotter et al. 2010, 2011), that is interpreted as a hint of an accreted origin. Specifically, due to the characteristics in the spaces defined by the integrals of motion, they have been assigned to GSE (NGC 362 and NGC 1261) and the Helmi Streams (Rup106). Also, we selected three targets (NGC 6218, NGC 6522 and NGC 6626) that are generally considered as in situ, since they fit the older branch of the AMR (Dotter et al. 2010, 2011, Villanova et al. 2017, Kerber et al. 2018) and move on a orbits compatible with either the MW disc (NGC 6218) or the Bulge (NGC 6522 and NGC 6626). The primary objective of this work is to test precise chemical tagging as a tool to infer the origin of GCs. When complementing the chrono-dynamical information, chemical tagging might clarify the differences among clusters with ambiguous associations. Furthermore, it enables determining if the progenitor galaxy and the associated GC candidates are chemically compatible. For this to be effective, the combination of high precision in abundances derivation, which can only be achieved using high-quality data, and homogeneity in the detailed analysis is required. This is why we reanalyze these six clusters, which have extensively been studied in the literature, to compare them with each other with the utmost precision.
@@
The paper is structured as follows. In Section 2 we list the observation that were collected to compile the spectroscopic dataset. In Section 3 we describe the methodology employed to carry out the chemical analysis. In Section 4 we show the output of the analysis. In Section 5 we compare our results with literature. In Section 6 we discuss the peculiarity of Ruprecht 106. In Section 7 we recap the main results of this work.
@@ 2 Observations
We retrieved our entire spectroscopic dataset from the ESO archive. The spectra of all the stars have been acquired using the multi-object spectrograph UVES-FLAMES (Pasquini et al. 2002) or UVES (Dekker et al. 2000) mounted at the Very Large Telescope (VLT) of the European Southern Observatory. Stars have been observed using the Red Arm 580 CD3 grating with a spectral coverage between 4800 and 6800 Å and a spectral resolution of R=40,000. Typical signal-to-noise ratios for these spectra are in the range of 40 S/N 65 at 5000 Å and 60 S/N 90 at 6000 Å. The spectra were reduced using the dedicated ESO pipeline111https://www.eso.org/sci/software/pipelines/ that includes bias subtraction, flat-fielding, wavelength calibration, spectral extraction and order merging. For the stars observed with UVES at VLT, the contribution of the sky has been properly removed from each stellar spectrum during the pipeline reduction. For the stars observed with UVES-FLAMES at VLT, the sky contribution have been taken into account by acquiring the spectra of some nearby sky regions at the same time as the science targets and subtracting it to each individual exposure. In the end, single exposures of the same star have been merged to obtain the final spectrum for each target. We display in Fig. 1 the spectra of six stars, one per target cluster, to showcase the quality of the dataset.
@@
As mentioned in the Introduction, the sample of GCs subject of this study has been selected based on their similar metallicity and on their dynamical association (always consistent among different studies, e.g., Massari et al. 2019, Forbes 2020, Callingham et al. 2022, see below). The individual spectroscopic targets analysed in this work has been selected as follows:
@@
NGC 362. This cluster has been categorized as accreted during the merger with GSE. The mean metallicity of NGC 362 is (Carretta et al. 2013). We collected the spectra of 14 RGB stars observed under the ESO-VLT Programme 083.D-0208 (PI: E. Carretta).
@@
NGC 1261. This GC has also been associated with the GSE accretion event. The mean metallicity of this cluster is (Marino et al. 2021). Our dataset includes the spectra of 12 RGB stars. Among them, 6 were observed under the ESO-VLT Programme 0101.D-0109 (PI: A. Marino), 4 were observed under the ESO-VLT Programme 197.B-1074 (PI: G. F. Gilmore) and 2 were observed under the ESO-VLT Programme 193.D-0232 (PI: F. R. Ferraro).
@@
NGC 6218. This system is an in-situ GC located in the MW disc. NGC 6218 has a mean metallicity of (Carretta et al. 2009). We include in our analysis 11 RGB stars observed under the ESO-VLT Programme 073.D-0211 (PI: E. Carretta).
@@
NGC 6522. This ia an in situ cluster placed in the MW Bulge. Its mean metallicity is (Barbuy et al. 2021). Our sample is composed of 4 RGB stars observed under the ESO-VLT Programme 097.D-0175 (PI: B. Barbuy).
@@
NGC 6626. This is also an in-situ MW Bulge globular cluster with a mean metallicity of (Villanova et al. 2017). We collected the spectra of 16 RGB stars observed under the ESO-VLT Programme 091.D-0535 (PI: C. Moni Bidin).
@@
Ruprecht 106. Among the population of accreted GCs, one that has always gathered great attention in the community is Rup106. Its orbit is typical of the Galactic halo (Frelijj et al. 2021) and due to its positioning in the spaces defined by the integrals of motion, this GC has been tentatively linked with the progenitor of the Helmi Streams. The chemical peculiarity of Rup106 has been revealed by several spectroscopic analyses (e.g., Brown et al. 1997, Villanova et al. 2013, V13 hereafter), which find that all Rup106 target spectra are depleted in [/Fe] compared to GCs and field stars of the same metallicity, that is (Lucertini et al. 2023). We collected the spectra of 9 RGB stars that have been observed under the ESO-VLT Programme 069.D-0642 (PI: P. François).
@@ 3.1 Stellar parameters
have been determined using the empirically calibrated - relation provided by Mucciarelli et al. (2021a). The choice of using the color is guided by the fact that it is the most extended in wavelength among Gaia colors, thus ensuring an optimal sampling of the spectral energy distribution. This approach also eliminates the need to rely on measurements from other photometric systems, which may have lower precision than Gaia. We checked that using other Gaia colours would result in differences always smaller than 50 K, which is below the current uncertainties. The color excess E(B-V) adopted for each GC is taken from Harris (2010). Given the high extinction in the direction of the Galactic bulge, we corrected the Gaia photometry of NGC 6522 and NGC 6626 for effects of differential reddening following the prescription described in Milone et al. (2012), using the Cardelli et al. (1989) exctinction law. This analysis was performed on the catalogues provided by Vasiliev & Baumgardt (2021), focusing only on stars with a probability membership . The effects of the extinction on the observed color have been taken into account following the iterative prescription described in Gaia Collaboration et al. (2018). Since the color - relation we employed is sensitive to the metallicity of the star, we assumed the value listed in the Harris (2010) catalogue as representative of the GC under investigation to get an initial estimate of . We note that estimating the using the derived values leads to variations from the initial values always lower than 20 K. Internal errors in due to the uncertainties in photometric data, reddening and assumed color - relation are in the range 80 - 110 K.
Surface gravities have been estimated from the Stefan-Boltzmann relation, using the photometric and a representative stellar mass of 0.8 , typical for RGB stars in old isochrones at this metallicity, the G-band bolometric corrections described by Andrae et al. (2018), and the distance provided by Baumgardt & Vasiliev (2021). The uncertainties on the log are derived through the propagation of the errors on , photometry and distance, and they are always under .
In the end, we calculated microturbulent velocities () by minimizing the trend between iron abundances and reduced equivalent widths, defined as (EW/). To do so, we used on average iron lines per star. We assumed a conservative uncertainty of 0.2 km . All the atmospheric parameters are listed in Table 1.

### v2

@@
Globular clusters (GCs) are extremely intriguing systems that help in reconstructing the assembly of the Milky Way via the characterisation of their chemo-chrono-dynamical properties. In this study, we use high-resolution spectroscopic archival data from UVES and UVES-FLAMES at the VLT to compare the chemistry of GCs dynamically tagged as either Galactic (NGC 6218, NGC 6522, and NGC 6626) or accreted from distinct merger events (NGC 362 and NGC 1261 from Gaia-Sausage-Enceladus, and Ruprecht 106 from the Helmi Streams) in the metallicity regime where abundance patterns of field stars with different origin effectively separate . We find remarkable similarities in the abundances of the two Gaia-Sausage-Enceladus GCs across all chemical elements. They both display depletion in the -elements (Mg, Si and Ca) and statistically significant differences in Zn and Eu compared to in situ GCs. Additionally, we confirm that Ruprecht 106 exhibits a completely different chemical makeup from the other target clusters, being underabundant in all chemical elements. This demonstrates that when high precision is achieved, the abundances of certain chemical elements can not only efficiently separate in situ from accreted GCs, but can also distinguish among GCs born in different progenitor galaxies. In the end, we investigate the possible origin of the chemical peculiarity of Ruprecht 106. Given that its abundances do not match the chemical patterns of the field stars associated with its most likely parent galaxy (i.e. the Helmi Streams), being depleted in the abundances of -elements in particular, we believe Ruprecht 106 to originate from a less massive galaxy compared to the progenitor of the Helmi Streams.
@@
The latest data releases of the ESA/Gaia mission (Gaia Collaboration et al. 2021, 2023) have instigated a profound transformation of our comprehension of the early chronicles of the Milky Way (MW). This enormous 6D phase space dataset has led to a clearer understanding of the roles played by various mergers in shaping the Galactic halo (see Helmi 2020 for a review), as predicted by simulations in the CDM cosmological framework (White & Frenk 1991, Moore et al. 1999, Helmi & de Zeeuw 2000, Newton et al. 2018). The imprint of accretion events manifests dynamically as stellar streams (Helmi et al. 1999, Ibata et al. 2024), substructures discernible in phase space (e.g. Helmi et al. 2018, Belokurov et al. 2018, Myeong et al. 2018, Koppelman et al. 2019), and dwarf galaxies whose disruption is currently ongoing (Ibata et al. 1994, Majewski et al. 2003). During such events, not only field stars but also globular clusters (GCs) can survive the merging processes, contributing to the formation of the present-day MW GC system (Brodie & Strader 2006, Peñarrubia et al. 2009, Bellazzini et al. 2020, Trujillo-Gomez et al. 2021). Leveraging the extremely precise kinematic data provided by the Gaia mission, the orbits of MW GCs have been reconstructed with exceptional precision. Through the analysis of their orbital properties, MW GCs have been classified into distinct groups by several authors (Massari et al. 2019, Forbes 2020, Callingham et al. 2022, Chen & Gnedin 2024), distinguishing those accreted from disrupted dwarf galaxies, such as Gaia-Sausage-Enceladus (GSE, Belokurov et al. 2018, Helmi et al. 2018), Sequoia (Myeong et al. 2018), and the Helmi Streams (Helmi et al. 1999), from those formed in situ. Complications may arise in the interpretation of these findings due to the complex and overlapping distribution of GCs in the spaces defined by their orbital parameters (Callingham et al. 2022). This complexity is further exacerbated when accounting for the impact of a non-static potential in the computation of a GC orbit (Amarante et al. 2022, Belokurov et al. 2023, Pagnini et al. 2023, Chen & Gnedin 2024). Nonetheless, it has been established that dynamically tagged groups of GCs lay on different sequences in the age–metallicity space (Kruijssen et al. 2019, Massari et al. 2019, Myeong et al. 2019), confirming the dual nature of the MW GC system in this plane already found in the pre-Gaia era (Marín-Franch et al. 2009, Forbes & Bridges 2010, Leaman et al. 2013, VandenBerg et al. 2013). In particular, it has been demonstrated that older systems at fixed metallicity are more likely to be born in galaxies with a higher star formation efficiency —such as the MW in its earliest phases— compared to dwarf galaxies (e.g. Massari et al. 2023).
@@
This intricate perspective of the assembly history of the MW can be further enriched by the addition of chemical abundance data obtained either from high-resolution spectra or large spectroscopic surveys (Venn et al. 2004, Nissen & Schuster 2010, Helmi et al. 2018, Horta et al. 2020, Minelli et al. 2021, Limberg et al. 2022, Malhan et al. 2022, Naidu et al. 2022a, Horta et al. 2023, Monty et al. 2023a, Ceccarelli et al. 2024). However, when looking at the distribution of GCs in chemical spaces, the differences between these two populations are subtle, complicating the task of interpreting the results. For instance, Recio-Blanco (2018) highlights a very low scatter in the ratio among GCs with regardless of their origin. On a similar note, Horta et al. (2020) employed the [Si/Fe] ratio from APOGEE DR17 (Abdurro’uf et al. 2022) of 46 GCs to explore whether their distribution in chemical spaces reflects the kinematically defined classification by Massari et al. (2019). Interestingly, Horta et al. (2020) observed that the positions of both in situ and accreted subgroups in this chemical space align with those of their field-star counterparts. However, they were unable to effectively identify different behaviours from GCs that were brought in by separate progenitors, possibly because of the higher precision needed. Recently, Belokurov & Kravtsov (2024) discriminated between Galactic and extragalactic origin for MW GCs using the [Al/Fe] ratios, building upon the findings of Belokurov & Kravtsov (2022). However, caution should be exercised when using the abundances of light elements in GCs owing to the impact of multiple populations (Bastian & Lardo 2018). Finally, Monty et al. (2023a, b) showcased the efficacy of differential chemical analysis on GCs. Their works not only revealed chemical inhomogeneity at a precision level of 0.02 dex for all chemical elements in GCs, but also showed distinctions in the chemical compositions of two GCs, NGC 288 and NGC 362, associated with the same progenitor, GSE. Specifically, NGC 288 and NGC 362 exhibit dissimilarities in their chemical makeup, particularly in their neutron-capture elements. This difference could be interpreted either as stemming from chemical inhomogeneities in the progenitor or as a consequence of internal chemical evolution within the GCs. In light of all the mentioned complexities, which underline the challenges associated with interpreting results when comparing diverse studies, it becomes evident that maintaining homogeneity is imperative in conducting chemical analyses aimed at verifying the origin of GCs.
@@
In the present study, we derived fully homogeneous detailed elemental abundances of several species for six GCs with similar metallicity that have been consistently associated with different progenitors (see below) according to all post-Gaia investigations (Massari et al. 2019, Forbes 2020, Callingham et al. 2022, Chen & Gnedin 2024). These GCs lie in the metallicity range best suited to chemical tagging studies, where abundance trends of in situ and accreted objects start to significantly part ways (Nissen & Schuster 2010, Helmi et al. 2018, Matsuno et al. 2022b). Three of the target clusters, namely NGC 362, NGC 1261, and Ruprecht 106 (Rup106 hereafter), lie on the younger branch of the Galactic GC age–metallicity relation (AMR, Dotter et al. 2010, 2011), which is interpreted as suggesting an accreted origin. Specifically, due to the characteristics of the spaces defined by the integrals of motion, they have been assigned to GSE (NGC 362 and NGC 1261) and the Helmi Streams (Rup106). Also, we selected three targets (NGC 6218, NGC 6522, and NGC 6626) that are generally considered to have formed in situ, as they fit the older branch of the AMR (Dotter et al. 2010, 2011, Villanova et al. 2017, Kerber et al. 2018) and move on orbits compatible with either the MW disc (NGC 6218) or the MW bulge (NGC 6522 and NGC 6626). The primary objective of the present work is to test precise chemical tagging as a tool to infer the origin of GCs. When complementing chrono-dynamical information, it is possible to use chemical tagging to clarify the differences among clusters with ambiguous associations. Furthermore, this approach enables determination of whether or not the progenitor galaxy and the associated GC candidates are chemically compatible. For this to be effective, we require a combination of high precision in abundance derivations, which can only be achieved using high-quality data, and homogeneity in the detailed analysis. This is why we reanalysed these six clusters, allowing us to compare them with each other with the utmost precision.
@@
The paper is structured as follows. In Sect. 2 we list the observations that were collected to compile the spectroscopic dataset. In Sect. 3 we describe the methodology employed to carry out the chemical analysis. In Sect. 4 we show the output of the analysis. In Sect. 5 we compare our results with the literature. In Sect. 6 we discuss the peculiarity of Ruprecht 106. In Sect. 7 we summarise the main results of this work.
@@ 2 Observations
We retrieved our entire spectroscopic dataset from the ESO archive. The spectra of these stars were all acquired using the multi-object spectrograph UVES-FLAMES (Pasquini et al. 2002) or UVES (Dekker et al. 2000) mounted at the Very Large Telescope (VLT) of the European Southern Observatory. Stars were observed using the Red Arm 580 CD3 grating with a spectral coverage between 4800 and 6800 Å and a spectral resolution of R=40,000. Typical signal-to-noise ratios for these spectra are in the range of 40 S/N 65 at 5000 Å and 60 S/N 90 at 6000 Å. The spectra were reduced using the dedicated ESO pipeline111https://www.eso.org/sci/software/pipelines/, which includes bias subtraction, flat-fielding, wavelength calibration, spectral extraction, and order merging. For the stars observed with UVES at the VLT, the contribution of the sky was properly removed from each stellar spectrum during the pipeline reduction. For the stars observed with UVES-FLAMES at the VLT, the sky contribution was taken into account by acquiring the spectra of some nearby sky regions at the same time as the science targets and subtracting it from each individual exposure. In the end, single exposures of the same star were merged to obtain the final spectrum for each target. Fig. 1 shows the spectra of six stars, one per target cluster, to showcase the quality of the dataset.
@@
As mentioned in Sect. 1, the sample of GCs studied here was selected based on their similar metallicity and dynamical association (which is consistent among the different studies; e.g. Massari et al. 2019, Forbes 2020, Callingham et al. 2022, see below). The individual spectroscopic targets analysed in this work were selected as follows:
@@
NGC 362. This cluster has been categorised as accreted during the merger with GSE. The mean metallicity of NGC 362 is (Carretta et al. 2013). We collected the spectra of 14 RGB stars observed under the ESO-VLT Programme 083.D-0208 (PI: E. Carretta).
@@
NGC 1261. This GC has also been associated with the GSE accretion event. The mean metallicity of this cluster is (Marino et al. 2021). Our dataset includes the spectra of 12 RGB stars. Among them, 6 were observed under the ESO-VLT Programme 0101.D-0109 (PI: A. Marino), 4 under the ESO-VLT Programme 197.B-1074 (PI: G. F. Gilmore), and 2 under the ESO-VLT Programme 193.D-0232 (PI: F. R. Ferraro).
@@
NGC 6218. This system is an in situ GC located in the MW disc. NGC 6218 has a mean metallicity of (Carretta et al. 2009). We include in our analysis 11 RGB stars observed under the ESO-VLT Programme 073.D-0211 (PI: E. Carretta).
@@
NGC 6522. This ia an in situ cluster placed in the MW bulge. Its mean metallicity is (Barbuy et al. 2021). Our sample is composed of four RGB stars observed under the ESO-VLT Programme 097.D-0175 (PI: B. Barbuy).
@@
NGC 6626. This is also an in situ MW bulge GC with a mean metallicity of (Villanova et al. 2017). We collected the spectra of 16 RGB stars observed under the ESO-VLT Programme 091.D-0535 (PI: C. Moni Bidin).
@@
Ruprecht 106. Among the population of accreted GCs, one that has always attracted significant attention within the community is Rup106. Its orbit is typical of the Galactic halo (Frelijj et al. 2021) and due to its positioning in the spaces defined by the integrals of motion, this GC has been tentatively linked with the progenitor of the Helmi Streams. The chemical peculiarity of Rup106 has been revealed by several spectroscopic analyses (e.g. Brown et al. 1997, Villanova et al. 2013, V13 hereafter), which find that all Rup106 target spectra are depleted in [/Fe] compared to GCs and field stars of the same metallicity, that is (Lucertini et al. 2023). We collected the spectra of nine RGB stars that have been observed under the ESO-VLT Programme 069.D-0642 (PI: P. François).
@@ 3.1 Stellar parameters
We determined the for all the targets of our sample using the empirically calibrated – relation provided by Mucciarelli et al. (2021a). The choice of using the colour is guided by the fact that it is the most extended in wavelength among Gaia colours, thus ensuring an optimal sampling of the spectral energy distribution. This approach also eliminates the need to rely on measurements from other photometric systems, which may have lower precision than Gaia. We verified that using other Gaia colours results in differences that are consistently smaller than 50 K, which is below the current uncertainties. The colour excess E(B-V) adopted for each GC is taken from Harris (2010). Given the high extinction in the direction of the Galactic bulge, we corrected the Gaia photometry of NGC 6522 and NGC 6626 for effects of differential reddening following the prescription described in Milone et al. (2012), and using the Cardelli et al. (1989) extinction law. This analysis was performed on the catalogues provided by Vasiliev & Baumgardt (2021), focusing on stars with a probability membership of . The effects of the extinction on the observed colour have been taken into account following the iterative prescription described in Gaia Collaboration et al. (2018). As the colour– relation we employed is sensitive to the metallicity of the star, we assumed the value listed in the Harris (2010) catalogue as representative of the GC under investigation to get an initial estimate of . We note that estimating the using the derived values leads to differences from the initial values consistently lower than 20 K. Internal errors in due to the uncertainties in photometric data, reddening, and the assumed colour– relation are in the range of 80 - 110 K. We estimated surface gravities from the Stefan-Boltzmann relation using the photometric and a representative stellar mass of 0.8 , typical for RGB stars in old isochrones at this metallicity, the G-band bolometric corrections described by Andrae et al. (2018), and the distance provided by Baumgardt & Vasiliev (2021). The uncertainties on the log were derived through the propagation of the errors on , photometry, and distance, and they are always under . In the end, we calculated microturbulent velocities () by minimising the trend between iron abundances and reduced equivalent widths, defined as (EW/). To do so, we used iron lines per star on average. We assumed a conservative uncertainty of 0.2 km . All the atmospheric parameters are listed in Table 1.